# Full-waveform Approximation of Finite-Sized Acoustic Apertures: Forward and Adjoint Wavefields

### v20

@@ 1. Introduction
In the Kirchhoff–Helmholtz formula, the integrand is generally expressed in terms of the acoustic pressure and its normal derivative over the bounding surface. However, since these two quantities are connected through a second homogeneous surface integral relation, the solution wavefield can be represented in terms of only one of them—either the pressure or its normal derivative. Consequently, using these surface integral formulas, boundary conditions are imposed on the chosen solution space, in addition to enforcing Cauchy conditions that satisfy causality. The choice of solution space and boundary conditions depends on the physics of the problem.
Correspondingly, we consider two cases: 1) A vibrating piston is mounted in a rigid baffle, where the normal derivative $\frac{\partial p}{\partial\bm{n}}$ of the pressure field vanishes everywhere on the baffle except on the acoustic aperture, which corresponds to the front face of the piston. In this configuration, a monopole (or singlet) source is defined in terms of the negative normal derivative of the pressure, $\frac{\partial p}{\partial\bm{n}}$, or equivalently, $\rho_{0}\frac{\partial\bm{u}}{\partial t}\cdot\bm{n}$ [20], where $\rho_{0}$ is the ambient density, $\bm{u}$ is the particle velocity vector, and $\bm{n}$ is the outward-pointing unit normal vector to the surface. This gives rise to the monopole integral representation, in which the Green’s function acts on a monopole source term restricted to the surface. 2) Under a soft-baffle assumption, the pressure $p$ is assumed to vanish on the baffle surface except on the acoustic aperture. In this case, the wavefield is represented using a surface-restricted dipole (or doublet [^1] The terms “singlet” and “doublet” are less common but refer to generalizations of the classical “monopole” and “dipole” sources, respectively.) source, given by $p\bm{n}$, where $p$ denotes the scalar pressure. The soft-baffle representation involves the action of the normal derivative of the Green’s function on $p$, giving rise to the dipole integral formula.
@@
In the case of the dipole integral formulation, the integrand involves the normal derivative of the Green’s function evaluated on the aperture surface. Analytically, this normal derivative depends on the obliquity factor, or equivalently, the solid angle, which quantifies the angle subtended by an elemental surface area as observed from a given point in the domain. Accordingly, in the dipole case, the extension operator maps a dipole source supported on a surface to a near-surface, vector-valued force source term, which is added to the equation of motion in the wave system.
Most importantly, building upon this extension operator, this work introduces a restriction operator that maps the wavefield—solutions of the wave equation—to boundary data corresponding to the pressure (or its normal derivative) measured at the receiver surface. We then show for the practical cases of Dirichlet-type boundary data (pressure) that the adjoint of a forward operator incorporating this trace mapping coincides—up to a constant factor—with a time-reversed variant of an interior-field formulation of the dipole surface integral formula, evaluated on the receiver surface. The corresponding time-reversed system serves as a back-projection operator that accounts for the analytic angular sensitivity of finite-size receivers. Given that modeling acoustic apertures as omnidirectional point sources may lack sufficient accuracy—particularly in high-frequency regimes relevant to biomedical applications—the proposed adjoint operator offers a compelling alternative for incorporation into iterative frameworks, such as error minimization algorithms [32, 35, 36, 37, 38, 58], to solve inverse problems involving receivers of finite size [59, 60].
Outline. Section 2 introduces the wave equation in the time domain and explains how a unique solution can be obtained using homogeneous Cauchy conditions, which establish a causal relationship between the solution wavefield and a radiation source. A primary solution, describing the propagated wavefield in terms of a radiation source with $d$-dimensional support (a volumetric region for $d=3$) in full space-time domain, is presented. Integral formulas expressing the wavefield in terms of a surface source are then derived. It will be shown how Dirichlet or Neumann boundary conditions can be imposed to obtain unique solutions for an overdetermined system arising from these integral formulas.
@@
In Section 5, building upon these integral formulations, we introduce the forward operator in the context of ultrasound tomography. In this section, we derive a restriction operator that maps the wavefield—approximated in free space—onto the Neumann or Dirichlet boundary data supported on subsurfaces corresponding to the finite-sized receivers. Each of these subsurfaces is assumed to lie on a corresponding infinite plane.
@@ 2. Wave Equation in the Time Domain
This section considers the propagation of acoustic waves from a real-valued space-and-time-varying source in an infinite, isotropic, and homogeneous medium in free space. Let $\bm{x}=\big[x^{1},\dots,x^{d}\big]^{T}$ denote a spatial position in $\mathbb{R}^{d}$ with $d\in\{2,3\}$ as the number of dimensions. The analysis presented here is performed for $d=3$ but holds for $d=2$ by replacing volumes with surfaces and surfaces with lines, and using the 2D Green’s function. (For instance, a line source in a 3D medium produces acoustic waves that propagate as cylindrical waves, equivalent to an omnidirectional point source in a 2D medium.)
@@ 5. Forward Problem Formulation in the Ultrasound Tomography Context
Building on the integral representations derived in the previous sections, we now formulate the forward problem. A general and representative example of this forward problem arises in ultrasound tomography, which can then be specialized to other applications such as therapeutic ultrasound optimization and photoacoustic tomography. In this study, we focus on the formulation of the forward and adjoint (or time-reversal) operators in this context, with the aim of embedding them into either an iterative inverse-problem framework or a single-step backprojection operator. Building on the derived and approximated formulas, the associated minimization problems can be constructed using first-order or second-order optimization methods [36, 37, 38, 39], or Neumann-series frameworks [32]. The exploration of the latter is left to further studies.
@@ 5.1. Problem Setting in Ultrasound Tomography
It was shown that the wavefield in a half space can be represented via an integral formulation involving the field and its normal derivative evaluated on a bounding surface $\partial\nu$, where $\partial\nu$ is taken to be an infinite plane.
@@
Each emitter–receiver pair produces a distinct time trace. Spherical or cylindrical acquisition geometries for $\partial\Omega$ are often realized by assembling multiple such transducers. Accordingly, in this setting the forward operator should be represented as one of the following boundary-to-boundary maps: Dirichlet-to-Neumann, Dirichlet-to-Dirichlet, Neumann-to-Neumann, or Neumann-to-Dirichlet. Among these, the Neumann-to-Dirichlet forward map is most frequently encountered in practical applications, and is the primary focus of analysis in this study. (We will also analyze the Neumann-to-Neumann map to preserve generality.)
@@
$\begin{split}\mathcal{A}_{r}:C_{0}^{\infty}\left(\mathbb{R}^{d}\times\mathbb{R}^{+}\right)&\rightarrow C_{0}^{\infty}\left(\partial\Omega\times[0,T]\right),\\ \mathcal{A}_{r}\left[\mathcal{S}_{e}\right]&=y_{r,e},\end{split}$ (39)
where $\mathcal{S}_{e}\in C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{+})$ denotes the regularized source term given by the right-hand side of the wave equations (35) and (38), with integration taken over the emitting surface $\partial\nu:=\partial\nu_{e}\cap\partial\Omega$. Although the source $\mathcal{S}_{e}$ is physically active only over a finite time interval, we formally extend its definition to all of space-time by assuming $\mathcal{S}_{e}(\bm{x},t)=0$ for all $t>T_{s}$. The forward operator $\mathcal{A}_{r}$ maps this source to the resulting pressure field measured on the boundary $\partial\Omega$ over a finite time, i.e., $y_{r}\in C_{0}^{\infty}\left(\partial\Omega\times[0,T]\right)$. (Note that the actual support of $y_{r}$ is confined to the disk-shaped subsurface represented by $\partial\nu_{r}\cap\partial\Omega$.)
Remark 3.
This provides the most general definition. For example, the forward operator associated with the acoustic process in photoacoustic tomography can be viewed as a composite operator, where $\mathcal{A}_{r}$ acts on an operator that maps the locally induced instantaneous source distribution to a time-varying source $s\in C_{0}^{\infty}(\Omega^{-})$.
@@ 5.2. Emission Operator
$\begin{split}\mathcal{Z}_{e}&:C_{0}^{\infty}\big(\partial\Omega\times[0,T]\big)\rightarrow C_{0}^{\infty}\big(\mathbb{R}^{d}\times\mathbb{R}^{+}\big),\\ \mathcal{Z}_{e}[z_{e}](\bm{x},t)&:=\begin{cases}-2\displaystyle\int_{\partial\nu_{e}\cap\partial\Omega}dS(\bm{x}_{s})\,\delta_{b}(\bm{x}-\bm{x}_{s})\,z_{e}(\bm{x}_{s},t),&\text{if }t\leq T,\\ 0,&\text{otherwise},\end{cases}\end{split}$ (42)
@@
From the regularized wave equation (29), at an arbitrary but fixed time $t\in[0,T]$, the near-surface field $\mathcal{Z}_{e}[z_{e}^{N}]$ is added to the div-grad field $\rho_{0}\nabla\cdot\left(\frac{1}{\rho_{0}}\nabla p_{e}\right)$. It follows that in the limit $b\rightarrow 0^{+}$, for each fixed time $t\in[0,T]$, there is a singularity in the field $\rho_{0}\nabla\cdot\left(\frac{1}{\rho_{0}}\nabla p_{e}\right)$ across $\partial\nu_{e}\cap\partial\Omega$. However, for positive $b$, when a regularized Dirac delta distribution—constructed as a tensor product of sinc functions in Cartesian coordinates—is used, the jump $2\frac{\partial p_{e}}{\partial\bm{n}_{e,-}}$ across the boundary does not induce a singularity. Instead, the singularity is smeared over a narrow region surrounding the boundary, with its magnitude decaying smoothly with increasing distance from $\partial\nu_{e}\cap\partial\Omega$.
@@
We define the free-space operators $\mathcal{G}_{e}^{N}$ and $\mathcal{G}_{e}^{D}$, acting at an arbitrary but fixed time $t\in\mathbb{R}^{+}$, on a field in free space as:
$\begin{split}\mathcal{G}_{e}^{N}:C_{0}^{\infty}\left(\mathbb{R}^{d}\times\mathbb{R}^{+}\right)&\rightarrow C_{0}^{\infty}\left(\mathbb{R}^{d}\times\mathbb{R}^{+}\right),\quad\mathcal{G}_{e}^{N}[f](\bm{x},t):=f(\bm{x},t),\\ \mathcal{G}_{e}^{D}:C_{0}^{\infty}\left(\mathbb{R}^{d}\times\mathbb{R}^{+}\right)&\rightarrow C_{0}^{\infty}\left(\mathbb{R}^{d}\times\mathbb{R}^{+}\right),\quad\mathcal{G}_{e}^{D}[f](\bm{x},t):=\frac{\partial f}{\partial\bm{n}_{e,-}}(\bm{x},t).\end{split}$ (43)
Here, $\mathcal{G}_{e}^{N}$ denotes the Neumann-type free-space operator, which acts as the identity map for all emitters $e$. The operator $\mathcal{G}_{e}^{D}$ denotes the Dirichlet-type free-space operator for emitter $e$, applying the inward normal derivative with respect to the infinite plane $\partial\nu_{e}$.
@@ 5.3. Reception Operator
We define the restriction operator that maps a field $f$ in free space-time domain to the measured data $y$ supported on $\partial\Omega$ over the finite time interval $[0,T]$. For each receiver $r$, this operator is given by
@@
Here, $\delta_{b,\partial\nu_{r}}(\bm{x}_{s}-\bm{x})$ denotes a regularized Dirac delta distribution supported on the (d1)(d{-}1)-dimensional infinite hyperplane $\partial\nu_{r}$. It is defined analogously to the standard Dirac delta distribution $\delta_{b}$ in $\mathbb{R}^{d}$, but with support restricted to the infinite plane (units: $\mathrm{m}^{1-d}$), and satisfies the sifting property
@@
Let $f\in C_{0}^{\infty}\big(\mathbb{R}^{d}\times[0,T]\big)$ be a smooth field, with smoothness dependent on the bandwidth parameter $b$. Then, on the receiver surface $\partial\nu_{r}\cap\partial\Omega$, in the limit $b\to 0^{+}$, $f$ can be recovered via the restriction operator acting on its outward normal derivative in a narrow region surrounding the receiver surface. Specifically, the field $f$ can be extracted on the surface $\partial\nu_{r}\cap\partial\Omega$ via
$f(\bm{x}_{s},t)=\mathcal{R}_{r}\Big[\frac{\partial f}{\partial\bm{n}_{r,+}}(\bm{x},t)\Big],\quad(\bm{x}_{s},t)\in\big(\partial\nu_{r}\cap\partial\Omega\big)\times[0,T],$ (50)
where $\bm{n}_{r,+}$ denotes the outward-pointing unit normal vector on $\partial\nu_{r}\cap\partial\Omega$, and $\mathcal{R}_{r}$ is defined in Eq. (48), with
$\delta_{b,\partial\nu_{r}}(\bm{x}_{s}-\bm{x}^{\prime})=b\,\delta_{b}(\bm{x}_{s}-\bm{x}^{\prime}).$
@@
At an arbitrary but fixed time $t\in[0,T]$, the test functions $h,f\in C_{0}^{\infty}\big(\mathbb{R}^{d}\times[0,T]\big)$ satisfy:
$\begin{split}\int_{\partial\nu_{r}}dS\,hf=\int_{\nu_{r}^{\pm}}d\bm{x}\,h\,\frac{\partial f}{\partial\bm{n}_{r,\pm}}+\int_{\nu_{r}^{\pm}}d\bm{x}\,\frac{\partial h}{\partial\bm{n}_{r,\pm}}f\end{split}$ (51)
where $\bm{n}_{r,+}$ denotes the outward unit normal vector to $\partial\nu_{r}\cap\partial\Omega$. We then choose the test function as
@@
$\frac{1}{b}\,\Big[f\big|_{\partial\nu_{r}^{+}}-f\big|_{\partial\nu_{r}^{-}}\Big](\bm{x}_{s},t)\;=\;\frac{\partial f}{\partial\bm{n}_{r,+}}(\bm{x}_{s},t),\quad(\bm{x}_{s},t)\in\big(\partial\nu_{r}\cap\partial\Omega\big)\times[0,T],$ (53)
which, in the limit $b\to 0^{+}$, indicates the emergence of a distributional singularity in $\tfrac{\partial f}{\partial\bm{n}_{r,+}}$ due to the sharp jump of $f$ across the boundary, leading to the symmetric relation
$\lim_{b\to 0^{+}}\,\frac{2}{b}\,f\big|_{\partial\nu_{r}}(\bm{x}_{s},t)\;=\;\frac{\partial f}{\partial\bm{n}_{r,+}}(\bm{x}_{s},t),\quad(\bm{x}_{s},t)\in\big(\partial\nu_{r}\cap\partial\Omega\big)\times[0,T],$ (54)
where we have used the continuity of $\frac{\partial f}{\partial\bm{n}_{r,+}}$ across the boundary.
@@
Substituting the expression for the extension operator $\mathcal{Z}_{e}$, introduced in Definition (4), into the restriction operator $\mathcal{R}_{r}$, introduced in Definition (9), and taking the limit $b\to 0^{+}$, we obtain:
$\begin{split}\lim_{b\to 0^{+}}\,\mathcal{R}_{r}\circ\mathcal{Z}_{e}\left[f(\bm{x}_{s},t)\right]&=\lim_{b\to 0^{+}}\ \frac{1}{2}\int_{\mathbb{R}^{d}}d\bm{x}\,\delta_{b,\partial\nu_{r}}(\bm{x}_{s}-\bm{x})\ \Big[2\int_{\partial\nu_{e}\cap\partial\Omega}dS(\bm{x}^{\prime}_{s})\,\delta_{b}(\bm{x}-\bm{x}^{\prime}_{s})\,f(\bm{x}^{\prime}_{s},t)\Big]\\ &=\lim_{b\to 0^{+}}\frac{1}{2}\int_{\mathbb{R}^{d}}d\bm{x}\,\delta_{b,\partial\nu_{r}}(\bm{x}_{s}-\bm{x})\frac{\partial f(\bm{x},t)}{\partial\bm{n}}=f(\bm{x}_{s},t),\end{split}$
for all $(\bm{x}_{s},t)\in(\partial\nu_{r}\cap\partial\Omega)\times[0,T]$. Note that, since the unit normal vector $\bm{n}$ is outward-pointing, the minus sign in the extension operator is dropped. Since $r=e$, we have $\partial\nu_{r}=\partial\nu_{e}$, and the above expression recovers $f$. This completes the proof.

### v21

@@ 1. Introduction
Within the Kirchhoff–Helmholtz integral representation, the acoustic pressure wavefield in a solution domain bounded by a surface is expressed as an integral involving a linear combination of the pressure field and its outward normal derivative evaluated on that surface. However, because these two boundary quantities are linked by an additional homogeneous surface integral relation in a complementary domain, the radiated wavefield can equivalently be represented using only a single boundary quantity—namely, either the pressure or its normal derivative. Consequently, using these surface integral formulas, boundary conditions are imposed on the chosen solution space, in addition to enforcing Cauchy conditions that satisfy causality. The choice of solution space and boundary conditions depends on the physics of the problem.
Correspondingly, we consider two cases: 1) A vibrating piston is mounted in a rigid baffle, where the normal derivative $\frac{\partial p}{\partial\bm{n}}$ of the pressure field vanishes everywhere on the baffle except on the acoustic aperture, which corresponds to the front face of the piston. In this configuration, a monopole (or singlet) source is defined in terms of the negative normal derivative of the pressure, $\frac{\partial p}{\partial\bm{n}}$, or equivalently, $\rho_{0}\frac{\partial\bm{u}}{\partial t}\cdot\bm{n}$ [20], where $\rho_{0}$ is the ambient density, $\bm{u}$ is the particle velocity vector, and $\bm{n}$ is the outward-pointing unit normal vector to the surface. This gives rise to the monopole integral representation, in which the Green’s function acts on a monopole source term restricted to the surface. 2) Under a soft-baffle assumption, the pressure $p$ is assumed to vanish on the baffle surface except on the acoustic aperture. In this case, the wavefield is represented using a surface-restricted dipole (or doublet [^1] In the mathematical literature, the corresponding “monopole” and “dipole” integral formulas are typically expressed as “single-layer” and “double-layer” potentials, respectively.) source, given by $p\bm{n}$, where $p$ denotes the scalar pressure. The soft-baffle representation involves the action of the normal derivative of the Green’s function on $p$, giving rise to the dipole integral formula.
@@
In the case of the dipole integral formulation, the integrand involves the normal derivative of the Green’s function evaluated on the aperture surface. Analytically, this normal derivative depends on the obliquity factor, or equivalently, the solid angle, which quantifies the angle subtended by an elemental surface area as observed from a given point in the domain. Accordingly, in the dipole case, the extension operator maps a vector-valued dipole source supported on a surface to a near-surface force source term, which is added to the equation of motion in the wave system.
Most importantly, building upon this extension operator, this work introduces a restriction operator that maps the wavefield—solutions of the wave equation—to boundary data corresponding to the pressure (or its normal derivative) measured at the receiver surface. We then show for the practical cases of Dirichlet-type boundary data (pressure) that the adjoint of a forward operator incorporating this trace mapping coincides—up to a constant factor—with a time-reversed variant of an interior-field formulation of the dipole surface integral formula, evaluated on the receiver surface. The corresponding time-reversed system acts as a back-projection operator that inherently accounts for the frequency-dependent and analytically characterized angular sensitivity of finite-size receivers. Given that modeling acoustic apertures as omnidirectional point sources may lack sufficient accuracy—particularly in high-frequency regimes relevant to biomedical applications—the proposed adjoint operator offers a compelling alternative for incorporation into iterative frameworks, such as error minimization algorithms [32, 35, 36, 37, 38, 58], to solve inverse problems involving receivers of finite size [59, 60].
Outline. Section 2 introduces the wave equation in the time domain and explains how a unique solution can be obtained using homogeneous Cauchy conditions, which establish a causal relationship between the solution wavefield and a radiation source. A primary solution, describing the propagated wavefield in terms of a radiation source with $d$-dimensional support (a volumetric region for $d=3$) in full space-time domain, is presented. Integral formulas expressing the wavefield in terms of a surface source are then derived. It will be demonstrated how imposing either Dirichlet or Neumann boundary conditions yields corresponding solutions to the overdetermined system arising from these integral representations.
@@
In Section 5, building upon these integral formulations, we introduce the forward operator in the context of ultrasound tomography (and also discuss the special case of photoacoustic tomography). In this section, we derive a restriction operator that maps the wavefield—approximated in free space—onto the Neumann or Dirichlet boundary data supported on subsurfaces corresponding to the finite-sized receivers. Each of these subsurfaces is assumed to lie on a corresponding infinite plane.
@@ 2. Wave Equation in the Time Domain
This section considers the propagation of acoustic waves from a real-valued space-and-time-varying source in an infinite, isotropic, and homogeneous medium in free space. Let $\bm{x}=\left[x^{1},\dots,x^{d}\right]^{T}$ denote a spatial position in $\mathbb{R}^{d}$ with $d\in\{2,3\}$ as the number of dimensions. The analysis presented here is performed for $d=3$ but holds for $d=2$ by replacing volumes with surfaces and surfaces with lines, and using the 2D Green’s function. (For instance, a line source in a 3D medium produces acoustic waves that propagate as cylindrical waves, equivalent to an omnidirectional point source in a 2D medium.)
@@ 4.2. System of Wave Equations in Terms of Regularized Source $\mathcal{S}$
Remark 2.
Although the analytic monopole and dipole integral formulas in Eqs. (22) and (24) are derived in a half-space to match the physical configuration, the regularized wave equation in Eq. (29) is extended to free space in order to accommodate the regularized source. From a numerical standpoint, the wavefield propagating into the complementary half-space is effectively suppressed by absorbing boundary conditions—such as Perfectly Matched Layers—embedded in the implementation of the wave system (30).
@@ 5. Forward Problem Formulation in the Ultrasound Tomography Context
Building on the integral representations derived in the previous sections, we now formulate the forward problem. A general and representative example of this forward problem arises in ultrasound tomography, which can then be specialized to other applications such as therapeutic ultrasound optimization and photoacoustic tomography. In this study, we focus on the formulation of the forward and adjoint (or time-reversal [^2] The notion of an adjoint operator frequently appears in the context of iterative minimization of an objective function; by contrast, the terms time-reversal, backprojection, or inverse operator all refer to the same concept and are more closely associated with single-step approaches, although embedding these latter operators within iterative minimization frameworks, such as Neumann-series iterations, has been reported [32]. ) operators in this context, with the aim of embedding them into either an iterative inverse-problem framework or a single-step backprojection operator. Building on these derived operators and their equivalent full-waveform approximations, the associated minimization problems can be solved using first-order or second-order optimization methods [36, 37, 38, 39], or within Neumann-series frameworks [32]. Exploration of the latter is left to future studies.
@@ 5.1. Problem Setting in Ultrasound Tomography
It was shown that the wavefield in a half space can be represented via an integral formulation involving the field or its normal derivative evaluated on a bounding surface $\partial\nu$, where $\partial\nu$ is taken to be an infinite plane [61].
@@
Each emitter–receiver pair produces a distinct time trace. Spherical or cylindrical acquisition geometries for $\partial\Omega$ are often realized by assembling multiple such transducers. Accordingly, in this setting, the forward operator can be represented as one of the following boundary-to-boundary maps: Dirichlet-to-Neumann, Dirichlet-to-Dirichlet, Neumann-to-Neumann, or Neumann-to-Dirichlet. Among these, the to-Dirichlet forward maps are most frequently encountered in practical applications and constitute the primary focus of analysis in this study. (We will also consider the to-Neumann maps to preserve generality.)
5.1.1. Forward Operator in Ultrasound Tomography
@@
$\begin{split}\mathcal{A}:C_{0}^{\infty}\left(\mathbb{R}^{d}\times\mathbb{R}^{+}\right)&\rightarrow C_{0}^{\infty}\left(\partial\Omega\times[0,T]\right),\\ \mathcal{A}_{r}\left[\mathcal{S}_{e}\right]&=y_{r,e},\end{split}$ (39)
where $\mathcal{S}_{e}\in C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{+})$ denotes the regularized source term given by the right-hand side of the wave equations (35) and (38), with integration taken over the emitting surface $\partial\nu:=\partial\nu_{e}\cap\partial\Omega$. Although the source $\mathcal{S}_{e}$ is physically active only over a finite time interval, we formally extend its definition to all of space-time by assuming $\mathcal{S}_{e}(\bm{x},t)=0$ for all $t>T_{s}$. The forward operator $\mathcal{A}_{r}$ maps the source to the resulting pressure field measured on the boundary $\partial\Omega$ over a finite time, i.e., $y_{r}\in C_{0}^{\infty}\left(\partial\Omega\times[0,T]\right)$. (Note that the actual support of $y_{r}$ is confined to the disk-shaped subsurface represented by $\partial\nu_{r}\cap\partial\Omega$.)
5.1.2. Forward Operator in Photoacoustic Tomography
The formulation above provides the most general definition of the operator. As a concrete example, consider the forward operator associated with acoustic wave propagation in photoacoustic tomography. In this setting, the overall forward map is realized as the composition $\mathcal{A}_{r}\circ\mathcal{Q}$. Here, $\mathcal{Q}$ denotes the operator that converts the locally induced instantaneous pressure distribution $p_{0}\in C_{0}^{\infty}(\Omega^{-})$ into a time-dependent source $s_{\mathcal{Q}}\in C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{+})$, i.e.,
$\begin{split}\mathcal{Q}:C_{0}^{\infty}(\Omega^{-})&\rightarrow C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{+}),\\ \mathcal{Q}[p_{0}]&=s_{\mathcal{Q}},\end{split}$ (40)
where
$s_{\mathcal{Q}}(\bm{x},t)=\frac{1}{c(\bm{x})^{2}}\,p_{0}(\bm{x})\,\frac{d}{dt}\delta(t).$ (41)
@@ 5.2. Emission Operator
$\mathcal{Z}_{e}$ $:C_{0}^{\infty}\big(\partial\Omega\times[0,T]\big)\rightarrow C_{0}^{\infty}\big(\mathbb{R}^{d}\times[0,T]\big),$ (44)
$\mathcal{Z}_{e}[z_{e}](\bm{x},t)$ $:=-2\displaystyle\int_{\partial\nu_{e}\cap\partial\Omega}dS(\bm{x}_{s})\,\delta_{b}(\bm{x}-\bm{x}_{s})\,z_{e}(\bm{x}_{s},t),$ (45)
@@
From the regularized wave equation (29), with a regularized source term defined by the right-hand side of Eq. (35), at an arbitrary but fixed time $t\in[0,T]$, the near-surface field $\mathcal{Z}_{e}[z_{e}^{N}]$ is added to the div-grad field $\rho_{0}\nabla\cdot\left(\frac{1}{\rho_{0}}\nabla p_{e}\right)$. It follows that in the limit $b\rightarrow 0^{+}$, for each fixed time $t\in[0,T]$, there is a singularity in the field $\rho_{0}\nabla\cdot\left(\frac{1}{\rho_{0}}\nabla p_{e}\right)$ across $\partial\nu_{e}\cap\partial\Omega$. However, for positive $b$, when a regularized Dirac delta distribution—constructed as a tensor product of sinc functions in Cartesian coordinates—is used, the jump $2\frac{\partial p_{e}}{\partial\bm{n}_{e,-}}$ across the boundary does not induce a singularity. Instead, the singularity is smeared over a narrow region surrounding the boundary, with its magnitude decaying smoothly with increasing distance from $\partial\nu_{e}\cap\partial\Omega$.
@@
We define the free-space operators $\mathcal{G}_{e}^{N}$ and $\mathcal{G}_{e}^{D}$, acting on a field in free space as:
$\begin{split}\mathcal{G}_{e}^{N}:C_{0}^{\infty}\left(\mathbb{R}^{d}\times[0,T]\right)&\rightarrow C_{0}^{\infty}\left(\mathbb{R}^{d}\times\mathbb{R}^{+}\right),\quad\mathcal{G}_{e}^{N}[f](\bm{x},t):=f(\bm{x},t),\\ \mathcal{G}_{e}^{D}:C_{0}^{\infty}\left(\mathbb{R}^{d}\times[0,T]\right)&\rightarrow C_{0}^{\infty}\left(\mathbb{R}^{d}\times\mathbb{R}^{+}\right),\quad\mathcal{G}_{e}^{D}[f](\bm{x},t):=\frac{\partial f}{\partial\bm{n}_{e,-}}(\bm{x},t).\end{split}$ (46)
Here, $\mathcal{G}_{e}^{N}$ denotes the Neumann-type free-space operator associated with emitter $e$, which extends the field to the full time by setting it to zero for all $t>T$. Likewise, $\mathcal{G}_{e}^{D}$ denotes the Dirichlet-type free-space operator for emitter $e$, which first applies the inward normal derivative with respect to the infinite plane $\partial\nu_{e}$ and subsequently extends the resulting field to all times.
@@ 5.3. Reception Operator
We define a restriction operator that, at each arbitrary but fixed time $t\in[0,T]$, maps a field $f$ in free space to the measured data $y$ supported on $\partial\Omega$. For each receiver $r$, this operator is given by
@@
Here, $\delta_{b,\partial\nu_{r}}(\bm{x}_{s}-\bm{x})$ denotes a regularized Dirac delta distribution supported on the (d1)(d{-}1)-dimensional infinite hyperplane $\partial\nu_{r}$. It is defined analogously to the standard Dirac delta distribution $\delta_{b}$ in $\mathbb{R}^{d}$, but with support restricted to the infinite plane (units: $\mathrm{m}^{1-d}$). It satisfies the sifting property
@@
Let $f\in C_{0}^{\infty}\big(\mathbb{R}^{d}\times[0,T]\big)$ be a field consisting of a smooth component plus a smeared jump across the receiver surface $\partial\nu_{r}\cap\partial\Omega$, where the jump decays as one moves away from the receiver surface and whose smoothness depends on the bandwidth parameter $b$. Then, on the receiver surface, in the limit $b\to 0^{+}$, the field $f$ can be recovered by applying the restriction operator to its outward normal derivative in a narrow region surrounding the receiver surface. Specifically,
$f(\bm{x}_{s},t)=\mathcal{R}_{r}\Big[\frac{\partial f}{\partial\bm{n}_{r,+}}(\bm{x},t)\Big],\qquad(\bm{x}_{s},t)\in\big(\partial\nu_{r}\cap\partial\Omega\big)\times[0,T],$ (53)
where $\bm{n}_{r,+}$ is the outward-pointing unit normal vector on $\partial\nu_{r}\cap\partial\Omega$, and $\mathcal{R}_{r}$ is defined in (51). Moreover,
$\delta_{b,\partial\nu_{r}}(\bm{x}_{s}-\bm{x}^{\prime})=b\,\delta_{b}(\bm{x}_{s}-\bm{x}^{\prime}),$
with $\delta_{b,\partial\nu_{r}}$ denoting the surface-localized mollifier of bandwidth $b$.
@@
At an arbitrary but fixed time $t\in[0,T]$, the functions $h,f\in C_{0}^{\infty}\big(\mathbb{R}^{d}\times[0,T]\big)$ satisfy:
$\begin{split}\int_{\partial\nu_{r}}dS\,hf=\int_{\nu_{r}^{\pm}}d\bm{x}\,h\,\frac{\partial f}{\partial\bm{n}_{r,\pm}}+\int_{\nu_{r}^{\pm}}d\bm{x}\,\frac{\partial h}{\partial\bm{n}_{r,\pm}}f,\end{split}$ (54)
where $\bm{n}_{r,+}$ and $\bm{n}_{r,-}$ denote the outward and inward unit normal vectors, respectively, on the surface $\partial\nu_{r}\cap\partial\Omega$. We then choose the test function as
@@
$\frac{1}{b}\,\Big[f\big|_{\partial\nu_{r}^{+}}-f\big|_{\partial\nu_{r}^{-}}\Big](\bm{x}_{s},t)\;=\;\int_{\mathbb{R}^{d}}d\bm{x}\ \delta_{b}(\bm{x}_{s}-\bm{x})\,\frac{\partial f}{\partial\bm{n}_{r,+}}(\bm{x},t),\quad(\bm{x}_{s},t)\in\big(\partial\nu_{r}\cap\partial\Omega\big)\times[0,T],$ (56)
which, in the limit $b\to 0^{+}$, indicates the emergence of a distributional singularity in $\tfrac{\partial f}{\partial\bm{n}_{r,+}}$ due to the sharp jump of $2f$ across the boundary, leading to the symmetric relation
$\lim_{b\to 0^{+}}\,\frac{2}{b}\,f\big|_{\partial\nu_{r}}(\bm{x}_{s},t)\;=\;\int_{\mathbb{R}^{d}}d\bm{x}\ \delta_{b}(\bm{x}_{s}-\bm{x})\,\frac{\partial f}{\partial\bm{n}_{r,+}}(\bm{x},t),\quad(\bm{x}_{s},t)\in\big(\partial\nu_{r}\cap\partial\Omega\big)\times[0,T].$ (57)
@@
At each arbitrary but fixed time $t\in[0,T]$, by substituting the expression for the extension operator $\mathcal{Z}_{e}$, introduced in Definition (4), into the restriction operator $\mathcal{R}_{r}$, introduced in Definition (9), and taking the limit $b\to 0^{+}$, we obtain:
$\begin{split}\lim_{b\to 0^{+}}\,\mathcal{R}_{r}\circ\mathcal{Z}_{e}\left[f(\bm{x}_{s},t)\right]&=\lim_{b\to 0^{+}}\ \frac{1}{2}\int_{\mathbb{R}^{d}}d\bm{x}\,\delta_{b,\partial\nu_{r}}(\bm{x}_{s}-\bm{x})\ \Big[2\int_{\partial\nu_{e}\cap\partial\Omega}dS(\bm{x}^{\prime}_{s})\,\delta_{b}(\bm{x}-\bm{x}^{\prime}_{s})\,f(\bm{x}^{\prime}_{s},t)\Big]\\ &=\lim_{b\to 0^{+}}\frac{1}{2}\int_{\mathbb{R}^{d}}d\bm{x}\,\delta_{b,\partial\nu_{r}}(\bm{x}_{s}-\bm{x})\frac{\partial f(\bm{x},t)}{\partial\bm{n}_{e,+}}=f(\bm{x}_{s},t),\end{split}$
for all $(\bm{x}_{s},t)\in(\partial\nu_{r}\cap\partial\Omega)\times[0,T]$. Note that, since the unit normal vector $\bm{n}_{e}$ is outward-pointing, the minus sign in the extension operator is dropped. Since $r=e$, we have $\partial\nu_{r}=\partial\nu_{e}$ and $\bm{n}_{r}=\bm{n}_{e}$, and the above expression recovers $f$. This completes the proof.